Kayla has 1.5 m (squared) of sheet metal to build a storage box for firewood.
a) what is the surface area of the metal , in square centimeters?
b) What are the dimensions for this square based prism box with maximum volume including a lid ?
c)If the box doesnt have a lid, what are the demensions of the square based prism. Round off to the nearest tenth of a centimeters.hint( make a table of possible boxes)
any help will be apreciated
for a) i think the area is 150 cm squared
for B according to my calculations the dimensions are 5 x 5 x 5 cm and the volume is 125 cm (cubed)
I dont know if my answers are correct and I am stuck on c)
any help will bw greatly appreciated . Pleas ehelp me
I need help desperately

Answer by stanbon(48568)

Kayla has 1.5 m (squared) of sheet metal to build a storage box for firewood.
a) what is the surface area of the metal , in square centimeters?
b) What are the dimensions for this square based prism box with maximum volume including a lid ?
c)If the box doesnt have a lid, what are the demensions of the square based prism. Round off to the nearest tenth of a centimeters.hint( make a table of possible boxes)
any help will be apreciated
c) Area of the sheet = 1.5*1.5 = 2.25 sq cm
The box has 5 faces of equal area:
Area of a face = 2.25/5 = 0.45 cm
A side of a face = sqrt(0.45) = 0.67 cm
Volume of the box is (0.67)^3 = 0.3019 cubic cm
=========================
Cheers,
Stan H.